### v1

@@ 4.5 Fine-tuning for the Observation Operator
To compare the atmospheric state with satellite observations, the observation operator is essential [19]. The simulated radiance is computed at each observation location using a profile of atmospheric variables. We fine-tune the pre-trained model by taking the satellite’s attitude parameters and scan position as conditions to construct an observation operator ^$\hat{\mathcal{H}}^{s}:\left(\mathbf{x}(t),\theta^{s}(t)\right)\mapsto\mathbf{y}^{s}(t)$, where $\theta^{s}(t)$ denotes the auxiliary observational information at time $t$. The loss function used to train ^$\hat{\mathcal{H}}^{s}$ is as follows:
@@
This study selected channels 5 to 10 of AMSU-A and 3 to 5 of MHS for training observation operators and conducting assimilation experiments. This selection was motivated by the fact that channels 5 to 10 of AMSU-A are sensitive to atmospheric temperature profiles ranging from 50 hPa to 1000 hPa, while channels 3 to 5 of MHS are responsive to the atmospheric water vapour in the troposphere [39]. The fine-tuning process of each observation operator costs about 13.3 hours to run on 8 A100-40GB GPUs. Please refer to the Supplementary Text for details regarding the configuration of the AMSU-A and MHS observation operators and their performance in Tables S5 and S6.
@@ 4.6 Fine-tuning for Data Assimilation
As illustrated in Figure 6(c), we fine-tune the pre-trained model to function as a DA model by incorporating the gradient of the 4DVar cost function with respect to the background field, denoted as $\nabla_{\mathbf{x}^{b}(t_{0})}\mathcal{J}^{\text{4DVar}}(\mathbf{x}^{b}(t_{0}))$. This approach follows the derivation presented in 4DVarFormer [14]. However, unlike 4DVarFormer, XiChen outputs the representation of analysis increments in latent space rather than the analysis increments themselves. This representation is added to the embedding of the background field $\mathbf{x}^{b}(t_{0})$ and subsequently decoded by the HNO blocks to generate the analysis field.
In our 4DVar cost function, the observation covariance matrix of prepbufr observations is denoted as $\mathbf{R}^{c}=\varepsilon^{c}(\varepsilon^{c})^{T}$. The observation operator error $\hat{\mathbf{\varepsilon}}^{s}$ is used to construct the error covariance matrix, $\mathbf{R}^{s}=\hat{\varepsilon}^{s}(\hat{\varepsilon}^{s})^{T}$, of the satellite observations. Since we only need to compute the gradient of the initial value of the 4DVar cost function with respect to $\mathbf{x}^{b}$, there is no need to estimate $\mathbf{B}$. This simplifies the process and circumvents a challenging issue typically encountered in the traditional 4DVar method [50]. In addition, unlike methods used in Aardvark [22], GraphDOP [23], FuXi-DA [13], and FengWu-Adas [32], we incorporate a quality control process into the computation of the 4DVar cost function by excluding observations whose distances from the background field exceed 5 times the standard deviation of the variable. This approach can prevent excessive adjustment of localized regions in the background field caused by large discrepancies between observations and the background. For a detailed depiction of the computational flow of the 4DVar cost function, please refer to Figure S3.
@@
We initially fine-tuned the pre-trained model using GDAS prepbufr observations to develop a prepbufr DA model, which costs about 45 hours to run on 8 A100-40GB GPUs. Subsequently, the prepbufr DA model was further fine-tuned into an AMSU-A DA model, employing the analysis field generated by assimilating GDAS prepbufr as the background field, which costs about 94 hours to run on 8 A100-40GB GPUs. Finally, the AMSU-A DA model was sequentially fine-tuned to assimilate MHS satellite observations, with the assimilating order of AMSU-A and MHS randomized, which costs about 222 hours to run on 8 A100-40GB GPUs. Detailed configurations for the assimilation model training are provided in the Supplementary Text.
@@ 4.7 Dual DA framework to operationalize XiChen as a cycling forecasting system
As shown in Figure 6(e), to circumvent the temporal latency inherent in AI-based 4DVar methods [24, 14] when deploying XiChen as a continuous forecasting system for acquiring medium-range prognostic results, we have architected a dual assimilation framework. Within this framework, we establish a configuration encompassing a 12-hour DAW and a 3-hour DAW. In the 12-hour DAW, we systematically assimilate observations retrospectively from the analysis time point at intervals of 0, 3, 6, and 9 hours to derive an accurate analysis field. This analysis field serves as the input for XiChen to generate a 12-hour forecast, subsequently yielding the background field for the initialization of the subsequent DAW cycle. Concurrently, within the 3-hour DAW configuration, we assimilate observations contemporaneous with the analysis time to generate the initial condition to initialize XiChen for the provision of medium-range forecast products. This dual DA framework facilitates the real-time acquisition of medium-range weather forecasting results without necessitating the temporal delay associated with the availability of all observations within the complete 12-hour DAW. In comparison to the IFS HRES, XiChen can deliver medium-range forecast products with a temporal advantage of several hours, thereby providing accurate weather forecasting products essential for governmental decision-making processes and people’s travel plans expeditiously.
@@ 4.8 Verification metrics
This study aims to assess our model’s performance concerning the tasks of assimilation and forecasting. Consequently, this study comprehensively evaluates a one-year DA cycle and a 10-day medium-range forecast. The DA cycle is executed at 12-hour intervals for assimilations, with a DAW of 12 hours. Specifically, the DA cycle is run for the year 2023 at 00:00 UTC and 12:00 UTC each day, which corresponds to the initialization times for the 10-day forecast conducted by the IFS HRES, Pangu-Weather [5], GraphCast [8], and FengWu [9]. To evaluate medium-range forecasts, we follow the WeatherBench [36] and DABench [33], selecting 50 initial fields at 336-hour intervals for the medium-range forecast experiments. The first initial field at 00:00 UTC is set for January 1, 2023, while the first initial field at 12:00 UTC is on January 8, 2023.
@@
We evaluate assimilate and forecast skill for a given variable, $\mathbf{x}_{v}$, using a latitude-weighted Root Mean Square Error (RMSE) [36] given by
@@
Bias
We also computed the bias for a given variable, $\mathbf{x}_{v}$
$\text{Bias}_{v}=\frac{1}{HW}\sum^{H}_{j}\sum^{W}_{k}L(i)(\hat{\mathbf{x}}_{v,i,j}-\mathbf{x}^{t}_{v,i,j}).$ (8)
The closer the Bias is to 0, the better the results are.
@@
To study skillful forecast lead times, we also calculated the latitude-weighted Anomaly Correlation Coefficient (ACC) [36] according to
@@
where $\mathbf{C}_{v,I,j}$ denotes the climatological mean for a given variable $v$ and the day-of-year containing the validity time at the ${i,j}$ grid point. It is calculated referring to GraphCast [8], FengWu [9], and DABench [33]. The hourly climatological fields are computed using ERA5 data between 2010 and 2021. The higher the ACC, the better the results.
@@
Availability of data and materials All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Information. The DABench dataset is available at the Baidu Drive. AMSU-A and MHS observations can be accessed here.
@@
Author contribution K.J.R., L.B., B.H.D., and W.X.W. designed the project. K.J.R., L.B., W.C.N., and B.H.D. managed and oversaw the project. W.X.W., S.M., and Y.L.Z. selected and preprocessed the observations. W.X.W., T.K.Y., and T.H. performed the model training and evaluation. W.X.W., T.H., and L.L.H. improved the model design. W.X.W., T.K.Y., X.Y.L., K.F.D., and T.H. analyze the experiment results. W.X.W., B.F., and W.C.N. wrote and revised the manuscript.

### v2

@@ 4.5 Fine-tuning for the Observation Operator
To compare the atmospheric state with satellite observations, the observation operator is essential [liang2023machine]. The simulated radiance is computed at each observation location using a profile of atmospheric variables. We fine-tune the pre-trained model by taking the satellite’s attitude parameters and scan position as conditions to construct an observation operator $\hat{\mathcal{H}}^{s}:\left(\mathbf{x}(t),\theta^{s}(t)\right)\mapsto\mathbf{y}^{s}(t)$, where $\theta^{s}(t)$ denotes the auxiliary observational information at time $t$. The loss function used to train $\hat{\mathcal{H}}^{s}$ is as follows:
@@
This study selected channels 5 to 10 of AMSU-A and 3 to 5 of MHS for training observation operators and conducting assimilation experiments. This selection was motivated by the fact that channels 5 to 10 of AMSU-A are sensitive to atmospheric temperature profiles ranging from 50 hPa to 1000 hPa, while channels 3 to 5 of MHS are responsive to the atmospheric water vapour in the troposphere [zou2017impacts]. The fine-tuning process of each observation operator costs about 13.3 hours to run on 8 A100-40GB GPUs. Please refer to the Supplementary Text for details regarding the configuration of the AMSU-A and MHS observation operators and their performance in Tables S7 and S8.
@@ 4.6 Fine-tuning for Data Assimilation
As illustrated in Figure 9(c), we fine-tune the pre-trained model to function as a DA model by incorporating the gradient of the 4DVar cost function with respect to the background field, denoted as $\nabla_{\mathbf{x}^{b}(t_{0})}\mathcal{J}^{\text{4DVar}}(\mathbf{x}^{b}(t_{0}))$. This approach follows the derivation presented in 4DVarFormer [wang2024accurate]. However, unlike 4DVarFormer, XiChen outputs the representation of analysis increments in latent space rather than the analysis increments themselves. This representation is added to the embedding of the background field $\mathbf{x}^{b}(t_{0})$ and subsequently decoded by the HNO blocks to generate the analysis field.
In our 4DVar cost function, the observation covariance matrix of prepbufr observations is denoted as $\mathbf{R}^{c}=\varepsilon^{c}(\varepsilon^{c})^{T}$. The observation operator error $\hat{\mathbf{\varepsilon}}^{s}$ is used to construct the error covariance matrix, $\mathbf{R}^{s}=\hat{\varepsilon}^{s}(\hat{\varepsilon}^{s})^{T}$, of the satellite observations. Since we only need to compute the gradient of the initial value of the 4DVar cost function with respect to $\mathbf{x}^{b}$, there is no need to estimate $\mathbf{B}$. This simplifies the process and circumvents a challenging issue typically encountered in the traditional 4DVar method [huang2024adaptive]. In addition, unlike methods used in Aardvark [allen2025end], GraphDOP [alexe2024graphdop], FuXi-DA [xu2024fuxida], and FengWu-Adas [chen2023fengwuadas], we incorporate a quality control process into the computation of the 4DVar cost function by excluding observations whose distances from the background field exceed 5 times the standard deviation of the variable. This approach can prevent excessive adjustment of localized regions in the background field caused by large discrepancies between observations and the background. For a detailed depiction of the computational flow of the 4DVar cost function, please refer to Figure S3.
@@
We initially fine-tuned the pre-trained model using GDAS prepbufr observations to develop a prepbufr DA model, which costs about 45 hours to run on 8 A100-40GB GPUs. Subsequently, the prepbufr DA model was further fine-tuned into an MHS DA model, ASCAT DA model, AMSU-A DA model, and SATWND DA model. Detailed configurations for the assimilation model training are provided in the Supplementary Text.
@@ 4.7 Dual DA framework to operationalize XiChen as a cycling forecasting system
As shown in Figure 9(e), to circumvent the temporal latency inherent in AI-based 4DVar methods [xiao2024fengwu4dvar, wang2024accurate] when deploying XiChen as a continuous forecasting system for acquiring medium-range prognostic results, we have architected a dual assimilation framework. Within this framework, we establish a configuration encompassing a 12-hour DAW and a 6-hour DAW. In the 12-hour DAW, we systematically assimilate observations retrospectively from the analysis time point at intervals of 0, 3, 6, and 9 hours to derive an accurate analysis field. This analysis field serves as the input for XiChen to generate a 12-hour forecast, subsequently yielding the background field for the initialization of the subsequent DAW cycle. Concurrently, within the 6-hour DAW configuration, we assimilate observations retrospectively from the analysis time point at intervals of 0 and 3 hours to generate the initial condition to initialize XiChen for the provision of medium-range forecast products. The 12-hour DAW employs 11 ensemble members, each derived from the previous cycle’s analysis field with added perturbations, processed through a 12-hour forecast. The perturbation method, similar to that used in Pangu [bi2023accurate], generates perturbations using 3 octaves of Perlin noise with scales of 0.2, 0.1, and 0.05, and the number of periods along each axis (longitude and latitude) is set to 1, 2, and 4, respectively. The weight of each member is calculated as follows:
$\mathcal{J}^{\text{3DVar}}_{O}(\mathbf{x}(t_{0}))$ $=\frac{1}{2}\|\mathbf{y}(t_{0})-\mathcal{H}\left(\mathbf{x}(t_{0})\right)\|^{2}_{\mathbf{R}^{-1}},$ (7)
$\omega_{idx}$ $=\frac{\exp(-J^{3DVar}_{O}(\mathbf{x}_{idx}(t_{0})))}{\sum^{Ens}_{idx^{\prime}=0}\exp(-J^{3DVar}_{O}(\mathbf{x}_{idx^{\prime}}(t_{0})))},$ (8)
where $\mathcal{J}^{\text{3DVar}}_{O}(\mathbf{x}(t_{0}))$ represents the observation term of the three-dimensional variational cost function, $Ens$ is the number of ensemble members, idxidx and $idx^{\prime}$ are the indices of the ensemble members, $\omega_{idx}$ represents the weight of the idxidx ensemble member. In this context, $\mathbf{y}(t_{0})$ and $\mathbf{x}(t_{0})$ are the GDAS prepbufr observations and the ensemble analysis member at the analysis time, respectively.
This dual DA framework facilitates the real-time acquisition of medium-range weather forecasting results without necessitating the temporal delay associated with the availability of all observations within the complete 12-hour DAW.
@@ 4.8 Verification metrics
This study aims to assess our model’s performance concerning the tasks of assimilation and forecasting. Consequently, this study comprehensively evaluates a one-year DA cycle and a 10-day medium-range forecast. The DA cycle is executed at 12-hour intervals for assimilations, with a DAW of 12 hours. Specifically, the DA cycle is run for the year 2023 at 00:00 UTC and 12:00 UTC each day, which corresponds to the initialization times for the 10-day forecast conducted by the IFS HRES, Pangu-Weather [bi2023accurate], GraphCast [lam2023learning], and FengWu [ailabchen2025fengwu]. To evaluate medium-range forecasts, we follow the WeatherBench [rasp2020weatherbench] and DABench [wang2024benchmark], selecting 50 initial fields at 336-hour intervals for the medium-range forecast experiments. The first initial field at 00:00 UTC is set for January 1, 2023, while the first initial field at 12:00 UTC is on January 8, 2023.
@@
We evaluate assimilate and forecast skill for a given variable, $\mathbf{x}_{v}$, using a latitude-weighted Root Mean Square Error (RMSE) [rasp2020weatherbench] given by
@@
To study skillful forecast lead times, we also calculated the latitude-weighted Anomaly Correlation Coefficient (ACC) [rasp2020weatherbench] according to
@@
where $\mathbf{C}_{v,I,j}$ denotes the climatological mean for a given variable $v$ and the day-of-year containing the validity time at the i,j{i,j} grid point. It is calculated referring to GraphCast [lam2023learning], FengWu [ailabchen2025fengwu], and DABench [wang2024benchmark]. The hourly climatological fields are computed using ERA5 data between 2010 and 2021. The higher the ACC, the better the results.
@@
Availability of data and materials All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Information. The DABench dataset is available at the Baidu Drive. The satellite observation MHS, ASCAT, AMSU-A, and SATWND observations can be accessed here.
@@
Author contribution K.J.R., L.B., B.H.D., H.Z.L., W.M.Z., J.Q.S., and W.X.W. designed the project. K.J.R., L.B., W.C.N., and B.H.D. managed and oversaw the project. W.X.W., S.M., and Y.L.Z. selected and preprocessed the observations. W.X.W., T.K.Y., and T.H. performed the model training and evaluation. W.X.W., T.H., and L.L.H. improved the model design. W.X.W., T.K.Y., X.Y.L., K.F.D., and T.H. analyze the experiment results. W.X.W., B.F., and W.C.N. wrote and revised the manuscript.